Dukeville Company Interior Design Dukeville is an interior-design firm owned and managed by Jason Duke, a reputed interior designer Mr. Duke charges his clients $400/hr for his consultation services and serves clients for 50 hours a week on average. A typical engagement is between 20 and 40 hours. Mr. Duke uses Garza & Co., a local firm, to bill his clients and track receivables balances. He pays a monthly fee of $3,000 to Garza & Co. Mr. Duke is satisfied with Garzas services, but is considering an alternative of hiring a part-time book-keeper He estimates that an in-house bookkeeper would need about 120 hours a month to prepare and track billings. The market rate for bookkeeping services is $15 per hour, and total cost (with employment taxes and other charges) is $20 per hour Mr. Duke would have to spend about 6 hours of his own time every month supervisingthe bookkeeper, an increase from the 4 hours he spends each month working with Garza In addition, Mr. Duke would need to rent a small storage space in the building for $150/mo if he used a temp bookkeeper Mr. Duke notes that he already has an extra desk and chair that the temp bookkeeper can use; he paid $2,000 for it and it has a 5-year estimated useful life Mr. Duke seeks your advice to compare the 2 alternatives (outsource to Garza or hire a PT bookkepper). What would your advice be? Show details of your work. Provide your response in two parts; the first part should be a quantitative comparison and the second part should discuss qualitative factors
Answer #1

Garza outsourcing cost : $ 3,000

Hire a bookkeeper

Relevant costs :salary : $20*120 = $2,400

Opportunity cost: Additional supervisory time of Mr Duke = (6-4) *$400 = $800

Rent cost = $150

Depreciation cost is a fixed cost and will not be taken into consideration during decision making.

Total cost = $ 2,400+$800+$150 = $3,350

This it is cheaper to get work outsourced.Also, it reduces the additional burden on Mr Duke and he can spend this time in servicing his clients.Also, by outsourcing the work Mr Duke is not required to rent any small storage place.The extra desk and chair can be put to some other use otherwise bookkeeper would have used them.